OpenStudy (anonymous):

anyone knows how can i solve this partial differential equation? (-\partial^2/\partial x^2-1/x \partial/\partial x-\partial^2/\partial y^2+a)u+i(\partial/\partial y)v=0

OpenStudy (anonymous):

Fix your tex, no one can read that giberish

OpenStudy (loser66):

agree with jack17. Use draw box below, please

OpenStudy (amistre64):

\[(-\frac{\partial^2}{\partial x^2}-\frac 1x \frac{\partial}{\partial x}-\frac{\partial^2}{\partial y^2}+a)u+i(\frac{\partial}{\partial y})v=0\]